one acetic acid solution is 70% water and another is 30% water. how many liters of each solution should be mixed to produce 20 liters of a solution that is 40% water?
Answer by oberobic(2304)

With 'solution' problems, you have to keep track of how much pure stuff you need.
This question is tricky because the question presents the complementary percentages.
Solution 1 is 70% water, so it is 30% pure acetic acid. We normally call that a 30% solution.
Solution 2 is 30% water, so it is 70% pure acetic acid. We normally call that a 70% solution.
The goal is 20 liters of a 60% acetic acid solution.
.
x = volume of 30% solution
20 - x = volume of 70% solution
.
.3x = amount of pure acetic acid
.7(20-x) = the other amount of pure acetic acid.
.
20 liters of 60% acetic acid will have .6*20 = 12 liters of pure acetic acid.
.
.3x + .7(20-x) = 12
.
.3x + 14 - .7x = 12
-.4x = -2
x = -2/-.4 = 5 liters of 30% solution
which means there is
15 liters of 70% solution.
.
checking to see if we have 12 L of pure acetic acid
.
.3*5 + .7*15 = 1.5 + 10.5 = 12
.
So, 5 liters of 30% acetic acid + 15 liters of 70% acetic acid results in 20 liters of 60% acetic acid solution.
.
Reviewing the question, we need to give the complementary answer in terms of percentage of water.
.
Answer:
5 liters of 70% water + 15 liters of 30% water = 20 liters of 40% water